# There is no Definable Grauert Direct Image Theorem

### v1

@@ 1. Introduction
Fix any o-minimal structure which expands $\mathbb{R}_{\rm an}$ or more generally an o-minimal structure such that its expansion by the locally analytic functions is still o-minimal.
@@ 2.1. Definability of sections of the Picard scheme
Note that the image of $\sigma_{\mathcal{L}}$ is contained in an open and closed subscheme of $\mathrm{Pic}(X/S)$ which is of finite type over $S$, given by fixing finitely many Hilbert polynomials, so the notion of definability makes sense here. The idea of the proof of Proposition 2.1 is simply to copy [Sim94, Sec. 5] and add the word definable when needed. We present a concise version of Simpson’s construction.
@@ 2.2. The character variety and definable local systems of rank one
As we assume that $H_{1}(Y,\mathbb{Z})=\pi_{1}^{\rm ab}(Y)$ is not a torsion group, then $\mathrm{Char}(Y)$ is of positive dimension. So we find a non-constant morphism $\mathbb{G}_{m}\to\mathrm{Char}(Y)$ of algebraic groups which gives rise by Proposition 2.6 to a definable locally free $f^{-1}\mathcal{O}_{\mathbb{G}_{m}}$ sheaf $\mathbb{L}$ of rank one on $X=Y\times{\mathbb{G}_{m}}$. Consider the associated definable line bundle $\mathcal{L}=\mathbb{L}\otimes_{f^{-1}\mathcal{O}_{\mathbb{G}_{m}}}\mathcal{O}_{X}$ and the associated analytic map to the Picard scheme $\sigma_{\mathcal{L}}\colon\mathbb{G}_{m}\to\mathrm{Pic}(Y)$.
@@
For $s\in\mathbb{G}_{m}$ with $|s|=1$ the local system $\mathbb{L}_{s}$ on $Y$ is unitary. Non-isomorphic unitary local systems of rank one give rise to non-isomorphic line bundles by Hodge theory. So the images by $\sigma_{\mathcal{L}}$ of these infinitely unitary $s$ are pairwise different, i.e. $\sigma_{\mathcal{L}}$ is not constant. ∎
If the definable sheaves $f_{*}\mathcal{L}(n)$ were coherent for all $n>0$ then $\sigma_{\mathcal{L}}$ would be definable by Proposition 2.1. But then $\sigma_{\mathcal{L}}$ would be algebraic by the definable Chow theorem [PS09, Cor. 4.5]. However, algebraic maps from a rational variety to an abelian variety are constant, see for example [Mil86, Cor. 3.9], so $\sigma_{\mathcal{L}}$ is constant contradicting Claim 2.7.

### v2

@@ 1. Introduction
Fix any o-minimal structure which expands $\mathbb{R}_{\rm an}$.
@@
In the theorem $f_{*}$ denotes the pushforward of sheaves on the definable sites relative to the fixed o-minimal structure.
@@ 2.1. Definability of sections of the Picard scheme
Note that the image of $\sigma_{\mathcal{L}}$ is contained in an open and closed subscheme of $\mathrm{Pic}(X/S)$ which is of finite type over $S$, given by fixing finitely many Hilbert polynomials, so the notion of definability makes sense here. The idea of the proof of Proposition 2.1 is simply to copy [Sim94, Sec. 5] and add the word definable when needed. We are going to present a concise version of Simpson’s construction.
@@ 2.2. The character variety and definable local systems of rank one
As we assume that $H_{1}(Y,\mathbb{Z})\cong\pi_{1}^{\rm ab}(Y)$ is not a torsion group, it follows that $\mathrm{Char}(Y)$ has positive dimension. So we find a monomorphism $\mathbb{G}_{m}\to\mathrm{Char}(Y)$ of algebraic groups which gives rise by Proposition 2.6 to a definable locally free $f^{-1}\mathcal{O}_{\mathbb{G}_{m}}$ sheaf $\mathbb{L}$ of rank one on $X=Y\times{\mathbb{G}_{m}}$. Consider the associated definable line bundle $\mathcal{L}=\mathbb{L}\otimes_{f^{-1}\mathcal{O}_{\mathbb{G}_{m}}}\mathcal{O}_{X}$ and the associated analytic map to the Picard scheme $\sigma_{\mathcal{L}}\colon\mathbb{G}_{m}\to\mathrm{Pic}(Y)$.
@@
For $s\in\mathbb{G}_{m}$ with $|s|=1$ the local system $\mathbb{L}_{s}$ on $Y$ is unitary. Non-isomorphic unitary local systems of rank one give rise to non-isomorphic line bundles by Hodge theory. So the images by $\sigma_{\mathcal{L}}$ of these infinitely many unitary $s$ are pairwise different, i.e. $\sigma_{\mathcal{L}}$ is not constant. ∎
Finally, if the definable sheaves $f_{*}\mathcal{L}(n)$ were coherent for all $n>0$ then $\sigma_{\mathcal{L}}$ would be definable by Proposition 2.1. But then $\sigma_{\mathcal{L}}$ would be algebraic by the definable Chow theorem [PS09, Cor. 4.5]. However, algebraic maps from a rational variety to an abelian variety are constant, see for example [Mil86, Cor. 3.9], so $\sigma_{\mathcal{L}}$ is constant contradicting Claim 2.7.